40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 15 17:48:22 2011, Jenne, Update, IOO, Mode Cleaner resonating again 
    Reply  Tue Feb 15 21:45:08 2011, rana, Update, IOO, The MC TRANS Story a.png
Message ID: 4304     Entry time: Tue Feb 15 21:45:08 2011     In reply to: 4303
Author: rana 
Type: Update 
Category: IOO 
Subject: The MC TRANS Story 

I forgot to elog that last night I touched up the MC2_TRANS QPD setup. I was perplexed by it always going out of alignment so I investigated.

I found that the fork clamp for the steering mirror for the QPD was not tightened. Shame. The beam diameter was equal to the aperture of the QPD and was clipping. Double shame.

I added a lens and tightened the mounts and centered the beam at ~9 PM yesterday. You can see in the attached trend that the measured power went up by ~10%.

Later, there's a big gap where Valera and Steve change out the PMC. You can see that the MC REFL voltage goes from 4.5 V to 5 V (10% increase in the power delivered to the MC).

There's essentially no change in the total transmission - this indicates that although the PMC transmission is now higher by ~10%, the matching to the IMC has been degraded by an equivalent fraction.

Needs some mode matching work.

Attachment 1: a.png  88 kB  | Hide | Hide all
a.png
ELOG V3.1.3-